Question Help

A physician with a practice is currently serving 280 patients. The physician would like to administer a survey to his patients to measure their satisfaction level with his practice. A random sample of 25 patients had an average satisfaction score of 8.1 on a scale of​ 1-10. The sample standard deviation was 1.21

a. Construct a​ 99% confidence interval to estimate the average satisfaction score for the​ physician's practice.( round to two decimals)

ps-EVERYONE keeps getting this wrong and its really frustrating because I cant get it either :/
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 8.1

sample standard deviation = s = 1.21

sample size = n = 25

Degrees of freedom = df = n - 1 = 24

At 99% confidence level the t is ,

t\alpha /2,df = t0.005,24= 2.797

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.797 * ( 1.21/ \sqrt 25)

= 0.68

The 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

8.1 - 0.68 < \mu < 8.1 + 0.68

7.42 < \mu < 8.78

( 7.91 , 8.29 )